OpenStudy (anonymous):

According to a recent survey, 78% of adults in the world are happy with their air quality. The margin of error was 0.03%. Approximately how many people were surveyed?

OpenStudy (anonymous):

What is the first step?

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

What do you think is the first step?

OpenStudy (anonymous):

change 78 into a percent or something.

OpenStudy (anonymous):

2) Margin of error is NOT 0.03%, it is 3% = 0.03

OpenStudy (anonymous):

1)*

OpenStudy (anonymous):

78 is already given as a percent... Look back to your last question for the margin of error formula. I'm assuming this is also a 95% confidence level.

OpenStudy (anonymous):

okay whats next?

OpenStudy (anonymous):

@SithsAndGiggles Is correct, 78% is already a percentage.

OpenStudy (anonymous):

0.03 = 1.96*sqrt [0.78*0.22/n] n = {1.96/0.03*sqrt [0.78*0.22]}^2 = 733 Do you understand?

OpenStudy (anonymous):

yeah but i put 733 in earlier and it was wrong.

OpenStudy (anonymous):

What are the selections? Because I got 733.

OpenStudy (anonymous):

there's none you just have to fill in the blank

OpenStudy (anonymous):

@Jesstho.-. how do we know the margin of error is .03 and not .0003? Maybe they mean .03%.

OpenStudy (anonymous):

Oh yes, you're right. @SithsAndGiggles I think it's the way he put the question, I got sort of confused.

OpenStudy (anonymous):

It could still be a typo, but I'd assume it's as written. \[.0003=1.96\sqrt\frac{.78(1-.78)}n\]and solve for \(n\).

OpenStudy (anonymous):

can you write like you did my other question so it can send me straight to the calculator?

OpenStudy (anonymous):

I don't think he could put it like that

OpenStudy (anonymous):

You mean send you a link to a calculator so I just give you the answer? Sure, why not. http://www.wolframalpha.com/input/?i=1.96*Sqrt%5B.78%281-.78%29%2Fn%5D%3D.0003

OpenStudy (anonymous):

Oh and @SithsAndGiggles That looks about correct

OpenStudy (anonymous):

K! Thanks!

OpenStudy (anonymous):

And in case it's as @Jesstho.-. thought: http://www.wolframalpha.com/input/?i=1.96*Sqrt%5B.78%281-.78%29%2Fn%5D%3D.03

OpenStudy (anonymous):

Ohh, it's 732 o-o this confused me a lot, great job @SithsAndGiggles